OpenStudy (anonymous):

Simplify the expression. sin^(2)x-1/cos-x

OpenStudy (anonymous):

-sin x cos x sin x -cos x

OpenStudy (anonymous):

answer choices

OpenStudy (gorv):

is it 1/cosx @pvs285

OpenStudy (anonymous):

no

OpenStudy (anonymous):

(sin^(2)x-1)/cos-x

OpenStudy (anonymous):

does that clear it up?

OpenStudy (anonymous):

@gorv help plz

OpenStudy (gorv):

is it cos(-x) ???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

whats the answer?

OpenStudy (gorv):

cos-x=cosx

OpenStudy (gorv):

\[\sin^{2}x-1=\cos^{2}x\]

OpenStudy (gorv):

\[\cos^{2}x/cosx\]

OpenStudy (gorv):

sorry sin^2x + cos^2x=1 sin^2x-1=-cos^2x

OpenStudy (gorv):

ans -cosx

OpenStudy (gorv):

@pvs285

OpenStudy (anonymous):

@gorv u there

OpenStudy (gorv):

yeah

OpenStudy (anonymous):

is it -cosx?

OpenStudy (gorv):

-cosx